OpenStudy (anonymous):

Solve each system by elimination. 2x-8y=-30 -2x-8y=-18

OpenStudy (insa):

add the both equations which gives u -16y= -48 y= -48/-16 y=3

OpenStudy (insa):

now subsitute y in any equation... i would go for 2nd which gives -2x-24= -18 -2x= 24-18 -2x= 6 x= 6/-2 x= -3 so y= 3 and x= -3